# A3: Android Agent Arena for Mobile GUI Agents

### v1

@@ Abstract
AI agents have become increasingly prevalent in recent years, driven by significant advancements in the field of large language models (LLMs). Mobile GUI agents, a subset of AI agents, are designed to autonomously perform tasks on mobile devices. While numerous studies have introduced agents, datasets, and benchmarks to advance mobile GUI agent research, many existing datasets focus on static frame evaluations and fail to provide a comprehensive platform for assessing performance on real-world, in-the-wild tasks. To address this gap, we present Android Agent Arena (A3), a novel evaluation platform. Unlike existing in-the-wild systems, A3 offers: (1) meaningful and practical tasks, such as real-time online information retrieval and operational instructions; (2) a larger, more flexible action space, enabling compatibility with agents trained on any dataset; and (3) automated business-level LLM-based evaluation process. A3 includes 21 widely used general third-party apps and 201 tasks representative of common user scenarios, providing a robust foundation for evaluating mobile GUI agents in real-world situations and a new autonomous evaluation process for less human labor and coding expertise. The project is available at https://yuxiangchai.github.io/Android-Agent-Arena/.
@@
Liang Liu vivo AI Lab                        Guozhi Wang vivo AI Lab                        Shuai Ren vivo AI Lab
Siyuan Huang Shanghai Jiao Tong University                        Hongsheng Lithanks: Corresponding Author MMLab @ CUHK
Figure 1: Overview of Android Agent Arena. A3 contains over 200 tasks from 21 widely used apps. Tasks are categorized into operation, single-frame query and multi-frame query based on the task goal. Tasks are also split into three difficulty levels based on the human operation steps. A3 also integrates two evaluation methods for different use cases.
@@ 1 Introduction
The significant advancement in the Large Language Model (LLM) area has boosted the development of AI agents, which are able to autonomously complete tasks following the instructions from users. Existing mobile AI assistants such as Siri, Xiao AI, and Bixby have demonstrated the potential of mobile agents to facilitate interactions between human users and mobile devices. However, those assistants are only effective in managing the routine tasks such as reporting weather condition and performing web searches due to the nature that they use APIs to perform task automation. To broaden the capability of AI agents, researchers have proposed GUI agents, which leverage the extended world knowledge and robust capabilities of multimodal large language models (MLLM) to effectively complete tasks on third-party general apps without the reliance on the APIs.
@@
Several recent works (Rawles et al., 2024a; Xing et al., 2024; Xu et al., 2024; Lee et al., 2024; Zhang et al., 2023b) have introduced dynamic evaluation platforms for Android GUI agents. While these efforts represent significant progress, they suffer from critical limitations that hinder their effectiveness as comprehensive evaluation benchmarks. For instance, many platforms restrict app selection to Google apps, F-Droid apps (non-mainstream open-source apps), or static offline apps, which do not reflect the diversity or complexity of real-world usage. Additionally, these platforms often provide only a limited diversity of tasks, or fail to include information query tasks, which are essential for evaluating practical agent performance. To address these shortcomings, we propose Android Agent Arena (A3), a novel evaluation system that offers: (i) integration with 21 widely used third-party apps and 201 tasks designed around real-world app functionalities, (ii) a diverse set of tasks categorized into three distinct types, and (iii) a larger action space, enabling compatibility with agents trained on any dataset. Furthermore, we introduce a new evaluation method that leverages the capabilities of business-level large language models (LLMs) to automate task evaluation, significantly reducing the need for human intervention and manual coding. Table 1 demonstrates the overview of A3.
@@
  • We introduce Android Agent Arena (A3), a comprehensive evaluation platform that integrates 201 tasks across 21 mainstream apps from real-world scenarios, with support for an expanded action space compatible with any dataset annotation style.
  • We propose a novel evaluation approach leveraging business-level LLMs, introducing two applications that enable scalable and automated evaluation processes while greatly reducing the need for manual effort in task evaluation.
Name Eval Mode # Tasks # General Apps Operation Inf. Query
@@
A3 (Our) dynamic 201 21
Table 1: GUI related datasets and benchmarks. 15 general apps of AndroidWorld are from F-Droid.
@@ 2.1 GUI Agent
(Wang et al., 2023) applies large language models (LLMs) to GUI tasks, but their focus remains limited to single-page interactions, resembling question-answering tasks rather than fully end-to-end instructional tasks. Recent advancements have begun leveraging the extensive world knowledge and robust reasoning capabilities of LLMs (Gu and Dao, 2024; Gur et al., 2023; Touvron et al., 2023) for task planning and execution. A notable approach involves using general-purpose business-level models, such as GPT-4v, as GUI-control agents. Studies like (Zhang et al., 2023a; Zheng et al., 2024) employ extensive prompt engineering to guide these models in performing complex tasks. However, the effectiveness of such methods depends heavily on meticulous prompt design to achieve satisfactory results. An alternative research direction focuses on fine-tuning smaller LLMs using GUI-specific datasets to imbue them with domain-specific knowledge, thereby improving their efficiency and task performance. For instance, CogAgent (Hong et al., 2024) enhances GUI task performance by incorporating a high-resolution cross-module that fuses image features at multiple levels. Similarly, MobileAgent (Ding, 2024) optimizes input data structuring and sample handling to make it more consistent and compatible with LLMs. SphAgent (Chai et al., 2024) utilizes the element functionalities to further enhance the screen and element understanding. CoCo-Agent (Ma et al., 2024) and AndroidControl (Li et al., 2024) take a different approach by integrating element layouts from accessibility trees or view hierarchies as additional inputs, rather than relying solely on screenshots. While this approach enhances performance, it faces limitations as many apps either lack accessibility information or provide only minimal accessibility data, which restricts its applicability in real-world scenarios.
@@ 2.3 Dynamic Evaluation Benchmark
To overcome the limitations of static frame evaluations, researchers have developed several dynamic evaluation systems (Lee et al., 2024; Xu et al., 2024; Xing et al., 2024; Rawles et al., 2024a; Zhang et al., 2023b) that aim to better simulate real-world environments (See Table 1). For instance, Mobile-Env (Zhang et al., 2023b) incorporates a broader range of general apps but is restricted to only 74 tasks. AndroidArena (Xing et al., 2024) introduces a larger number of tasks, including cross-app tasks, but is limited to Google apps and built-in system apps (e.g., settings and clock), which are already manageable by API-based assistants. B-Moca (Lee et al., 2024) supports a Korean language setting but offers tasks that are overly simplistic and lack diversity. AndroidWorld (Rawles et al., 2024a) uses open-source apps from F-Droid [^1] https://f-droid.org/en/packages/, but these apps differ significantly from mainstream app designs, making them unrepresentative of real-world scenarios. Notably, all of these systems focus only on operational instructions and their corresponding evaluations. AndroidLab (Xu et al., 2024) is the first system to incorporate information query instructions and evaluations, addressing a key gap. However, its app selection is limited to offline and static apps, failing to include categories such as news, shopping, email, and music, which are critical for real-world usability. To specify, the evaluation methods in existing systems predominantly rely on element matching (Lee et al., 2024) or predefined answers (Xu et al., 2024).
@@ 3.1 Overview
A3 is a lightweight system built on Appium [^2] https://appium.io/docs/en/latest/, an open-source framework for controlling Android and iOS devices. As shown in Figure 2, A3 acts as a bridge between a GUI agent and an Android device. It integrates tasks with their corresponding evaluation functions. The process begins with the controller retrieving the device’s current state, which includes a screenshot and an Extensible Markup Language (XML) file. This state and task instruction, along with additional information such as previous screenshots, XML files, and actions, is then sent to the agent. The agent analyzes the input and predicts the next action to take based on the current state. The predicted action is passed to the translator, which converts it into device control commands to interact with the device. This loop continues until the agent signals task completion or the predefined maximum number of steps is reached. At the end of the process, the evaluator determines whether the task was successfully completed using the evaluation function. The system is designed with flexibility and extensibility, allowing users to easily add new apps and tasks, and it also provides a universal translator system for any agent.
@@ 3.3 Task
Figure 3: Distribution of tasks in A3. The above subfigure is the distribution by categories and the bottom subfigure is the distribution by difficulty levels.
Unlike existing approaches, A3 incorporates over 200 tasks derived from 21 widely used third-party applications, thereby significantly broadening the scope and variety of real-world scenarios. Each task is deliberately chosen to represent the most common functionalities and use cases of a given application. Moreover, every task is distinct, minimizing the repetition of actions and intentions. To better characterize the types of tasks included, we classify them into three categories: (i) operation tasks, (ii) single-frame query tasks, and (iii) multi-frame query tasks.
@@
We divide all tasks into three tiers of difficulty. Tasks that a human can complete in four or fewer steps are considered easy, and those achievable in eight or fewer steps are deemed medium. All remaining operation tasks are classified as hard. Figure 3 illustrates the distribution of tasks in A3. Due to the fact that multi-frame query tasks are extremely hard for existing agents, we devote more effort in operation and single-frame query tasks for better evaluation of agents’ capabilities. Moreover, task instructions related to dates are dynamically generated to ensure the date can be selected, such as in a booking app.
@@ 3.4 Evaluation
In A3, we present two evaluation methods: (i) a task-specific evaluation function and (ii) a business-level LLM evaluation system. These methods operate independently and can be chosen by users, with the first focusing on predefined tasks and the second offering a scalable solution for adding tasks across various apps. And to mimic the real-world scenario, all tasks are evaluated by real-time states, which means all the contents are real-time, not predefined offline contents.
@@ 3.4.1 Evaluation Function
  • Element matching is the most commonly used evaluation method. It involves identifying key elements in an XML tree and comparing their attributes with ground-truth values. For instance, consider the task: "Open Downloads in Coursera and tell me how much storage is used." In this case, the final state XML should contain an element that displays the total storage used by the app. The ground-truth value can be extracted by parsing the XML tree and then compared to the agent’s response. In more complex scenarios, multiple elements may need to be retrieved, and several conditions must be met to determine if the agent’s answer is correct. Additionally, when the XML data is insufficient, OCR (Optical Character Recognition) is used to extract text attributes directly from an element, serving as a substitute for XML parsing.
  • Action matching is used when evaluation requires verifying specific positions. For example, in the task: "Search for a flashlight on the Wish app, filter results by price under 100, then select the first item and add it to the wishlist," the agent must click on the first item displayed in the search results. Action matching ensures that the click coordinates fall within the bounding box of the first item.
In more complex scenarios, both element matching and action matching are combined to accurately evaluate the agent’s performance.
@@ 3.4.2 LLM Evaluation System
LLM Correct Func. Wrong Line
GPT-4o 24% 27%
Table 2: The capability of GPT-4o to directly generate evaluation function. “Correct Func.” represents the percentage of correct files over all generated files. “Wrong Line” represents the percentage of incorrect lines over all code generated. The evaluation is collected from coding experts.
@@
One of the challenges faced by previous online evaluation systems is the difficulty of scaling tasks alongside their corresponding evaluation functions. During the development of A3, we also encountered this issue. Each task requires a specific evaluation function, which must be crafted by coding experts capable of parsing XML and defining precise conditions for task success. This dependency on manual coding hinders the rapid creation of evaluation functions. To address this limitation, we propose a business-level LLM evaluation system that leverages the advanced capabilities of large language models (LLMs) such as GPT and Gemini to enable semi-autonomous or even fully autonomous task evaluation.
As a first step, we utilize GPT-4o’s coding abilities to generate evaluation functions for tasks. Specifically, we supply GPT with our XML parsing code, an example evaluation function, and task-specific prompts to generate the required evaluation functions. While GPT-4o demonstrates strong coding capabilities, it can occasionally produce logic errors or incorrect conditions. For instance, when evaluating hotel searches on Booking.com, the search results dynamically change based on the selected date. However, GPT-4o might incorrectly attempt to match a static value, such as "113 results," instead of using a regex pattern containing the word "results" to identify whether the searching process is finished. To quantify GPT-4o’s performance, Table 2 illustrates its impact on task evaluation. Although the proportion of fully correct functions is relatively low, the percentage of lines requiring human modification is also minimal. This suggests that GPT-4o significantly reduces the coding workload, enabling a more efficient and semi-autonomous evaluation process.
The first evaluation method delivers accurate results due to human validation of the evaluation functions. However, it depends on coding experts, which can be a constraint. To overcome this, we propose a fully LLM-based evaluation process. For operation and single-frame query tasks, GPT-4o and Gemini 1.5 Pro are provided with task instructions and the corresponding final state XMLs, enabling them to directly assess task completion. For multi-frame query tasks, a sequence of XMLs is used to evaluate their ability to handle more complex scenarios. Table 3 highlights the evaluation accuracy of both LLMs, demonstrating approximately 80% correctness. To improve reliability, we introduce a cross-validation process that combines the outputs of both LLMs. When the two models produce the same result, the probability of misjudgment decreases to around 0.03, ensuring high confidence. In cases of disagreement, human evaluation is applied. This approach significantly reduces the reliance on human labor compared to the first method while also eliminating the need for coding expertise.

### v2

@@ Abstract
AI agents have become increasingly prevalent in recent years, driven by significant advancements in the field of large language models (LLMs). Mobile GUI agents, a subset of AI agents, are designed to autonomously perform tasks on mobile devices. While numerous studies have introduced agents, datasets, and benchmarks to advance mobile GUI agent research, many existing datasets focus on static frame evaluations and fail to provide a comprehensive platform for assessing performance on real-world, in-the-wild tasks. To address this gap, we present Android Agent Arena (A3), a novel evaluation platform. Unlike existing in-the-wild systems, A3 offers: (1) meaningful and practical tasks, including real-time online information retrieval and operational instructions; (2) a larger, more flexible action space, enabling compatibility with agents trained on any dataset; and (3) automated commercial LLM-based evaluation process which supports task customization. A3 includes 20 widely used general third-party apps and 201 tasks representative of common user scenarios, providing a robust foundation for evaluating mobile GUI agents in real-world situations and a new autonomous evaluation process for less human labor and coding expertise.
@@
Liang Liu vivo AI Lab                        Guangyi Liu vivo AI Lab                        Guozhi Wang vivo AI Lab
Shuai Ren vivo AI Lab                        Siyuan Huang Shanghai Jiao Tong University                        Hongsheng Lithanks: Corresponding Author MMLab @ CUHK
Figure 1: Overview of Android Agent Arena. A3 contains 201 tasks from 20 widely used apps. Tasks are categorized into operation, single-frame query and multi-frame query based on the task goal. Tasks are also split into three difficulty levels based on the human operation steps. A3 also integrates two evaluation methods for different use cases.
@@ 1 Introduction
Existing mobile AI assistants such as Siri, Xiao AI, and Bixby have demonstrated the potential of mobile agents to facilitate interactions between human users and mobile devices. However, those assistants are only effective in managing the routine tasks such as reporting weather condition and performing web searches due to the nature that they use APIs to perform task automation. To broaden the capability of AI agents, researchers have proposed GUI agents (Liu et al., 2025), which leverage the extended world knowledge and robust capabilities of multimodal large language models (MLLM) to effectively complete tasks on third-party general apps without the reliance on the APIs.
@@
Several recent works (Rawles et al., 2024a; Xing et al., 2024; Xu et al., 2024; Lee et al., 2024; Zhang et al., 2023b) have introduced dynamic evaluation platforms for Android GUI agents. While these efforts represent significant progress, they suffer from critical limitations that hinder their effectiveness as comprehensive evaluation benchmarks. For instance, many platforms restrict app selection to Google apps, F-Droid apps (non-mainstream open-source apps), or static offline apps, which do not reflect the diversity or complexity of real-world usage. Additionally, these platforms often provide only a limited diversity of tasks and do not support task customization, or fail to include information query tasks, which are essential for evaluating practical agent performance. To address these shortcomings, we propose Android Agent Arena (A3), a novel evaluation system that offers: (i) integration with 20 widely used third-party apps and 201 tasks designed around real-world app functionalities, (ii) a diverse set of tasks categorized into three distinct types, and (iii) a larger action space, enabling compatibility with agents trained on any dataset. Furthermore, we introduce a new evaluation method that leverages the capabilities of commercial LLMs to automate task evaluation, significantly reducing the need for human intervention and manual coding. Table 1 demonstrates the overview of A3.
@@
  • We introduce Android Agent Arena (A3), a comprehensive evaluation platform that integrates 201 diverse multi-category tasks across 20 mainstream apps from real-world scenarios, with support for an expanded action space compatible with any dataset annotation style.
  • We propose a novel evaluation approach leveraging commercial LLMs, which enables scalable and automated evaluation processes while greatly reducing the need for manual effort in task evaluation and keeping the evaluation precision, suitable for customized tasks evaluation.
Name Eval Mode # Tasks # General Apps Operation Inf. Query Online
@@
A3 (Our) dynamic 201 20
Table 1: GUI related datasets and benchmarks. The top three rows are GUI-related datasets, which provide static frame evaluation. The middle five rows are dynamic evaluation systems, which provide different tasks from different apps in different settings. AndroidWorld provides 15 generals apps from non-mainstream open-source F-Droid.
@@ 2.1 GUI Agent
Recent advancements have begun leveraging the extensive world knowledge and robust reasoning capabilities of LLMs (Gu and Dao, 2024; Gur et al., 2023; Touvron et al., 2023) for task planning and execution. A notable approach involves using general-purpose business-level models, such as GPT-4v, as GUI-control agents. Studies like (Zhang et al., 2023a; Zheng et al., 2024) employ extensive prompt engineering to guide these models in performing complex tasks. An alternative research direction focuses on fine-tuning smaller LLMs using GUI-specific datasets to imbue them with domain-specific knowledge, thereby improving their efficiency and task performance. For instance, CogAgent (Hong et al., 2024) enhances GUI task performance by incorporating a high-resolution cross-module that fuses image features at multiple levels. SphAgent (Chai et al., 2024) utilizes the element functionalities to further enhance the screen and element understanding. UI-TARS (Qin et al., 2025) finetunes the model with more reasoning data and GUI perception data. Qwen2.5-VL is a generalist Multimodal Large Language Model (MLLM) which pretrains on GUI-related data and achieved sound performance on various benchmarks.
@@ 2.3 Dynamic Evaluation Benchmark
To overcome the limitations of static frame evaluations, researchers have developed several dynamic evaluation systems (Lee et al., 2024; Xu et al., 2024; Xing et al., 2024; Rawles et al., 2024a; Zhang et al., 2023b) that aim to better simulate real-world environments (See Table 1). However, these systems still suffer from several shortcomings that limit their applicability and realism. For instance, Mobile-Env (Zhang et al., 2023b) is restricted to only 74 tasks, limiting its comprehensiveness of task diversity. AndroidArena (Xing et al., 2024) expands the task set by including cross-app interactions, yet it is restricted to Google apps and built-in system apps (e.g., settings and clock), which are already manageable by API-based assistants, thus failing to assess general usability across diverse third-party applications. B-Moca (Lee et al., 2024) introduces a Korean language setting, but its tasks are overly simplistic and lack diversity, making it inadequate for evaluating complex interactions (e.g., “input 1 in Calculator”, “go to account tab in Walmart”). AndroidWorld (Rawles et al., 2024a) attempts to introduce variety by utilizing open-source apps from F-Droid [^1] https://f-droid.org/en/packages/. However, these apps often differ significantly from mainstream app designs, making them unrepresentative of real-world usage scenarios. A critical limitation across all these systems is their narrow focus on operational instructions and corresponding evaluations, neglecting more complex real-world interactions. AndroidLab (Xu et al., 2024) makes an important step forward by incorporating information query instructions and evaluations, but it remains constrained by its reliance on offline and static apps. This limitation excludes crucial categories such as news, shopping, email, and music, which are integral to real-world user experiences. Additionally, the evaluation methodologies used in existing systems are often simplistic, primarily relying on element matching (Lee et al., 2024) or predefined answers (Xu et al., 2024), which fail to capture the nuances of naturalistic interactions and adaptive responses. Furthermore, all existing systems do not provide a solution for customized tasks evaluation, which is important for other common tasks not covered in the systems.
@@ 3.1 Overview
A3 is a lightweight system built on Appium [^2] https://appium.io/docs/en/latest/, an open-source framework for controlling Android and iOS devices. As shown in Figure 2, A3 acts as a bridge between a GUI agent and an Android device. It integrates tasks with their corresponding evaluation functions. The process begins with the controller retrieving the device’s current state, which includes a screenshot and an Extensible Markup Language (XML) file. This state and task instruction, along with additional information such as previous screenshots, XML files, and actions, is then sent to the agent. The agent analyzes the input and predicts the next action to take based on the current state. The predicted action is passed to the translator, which converts it into device control commands to interact with the device. This loop continues until the agent signals task completion or the predefined maximum number of steps is reached. For evaluation process, users can choose from function evaluation and commercial LLM evaluation, based on their needs and budgets. A3 also provides optional essential states evaluation for commercial LLM evaluation method to fully assess the ability of agents. The details of app selection and task generation is in Appendix A.1 and  A.2.
@@ 3.2 Action Space
Easy Med. Hard Op. S.Q M.Q
94 77 30 143 49 9
Table 2: The distribution of tasks in A3. “Op” stands for operation. “S.Q.” and “M.Q” stands for single-page query and multi-page query respectively.
@@ 3.3 Task
Unlike existing approaches, A3 incorporates over 200 tasks derived from 20 widely used third-party applications, thereby significantly broadening the scope and variety of real-world scenarios. Each task is deliberately chosen to represent the most common functionalities and use cases of a given application. Moreover, every task is distinct, minimizing the repetition of actions and intentions. To better characterize the types of tasks included, we classify them into three categories: (i) operation tasks, (ii) single-frame query tasks, and (iii) multi-frame query tasks.
@@
We divide all tasks into three tiers of difficulty. Tasks that a human can complete in five or fewer steps are considered easy, and those achievable in ten or fewer steps are deemed medium. All remaining operation tasks are classified as hard. Table 2 illustrates the distribution of tasks in A3. Due to the fact that multi-frame query tasks are extremely hard for existing agents, we devote more effort in operation and single-frame query tasks for better evaluation of agents’ capabilities. Moreover, task instructions related to dates are dynamically generated to ensure the date can be selected, such as in a hotel booking app.
@@ 3.4 Evaluation
In A3, we present two evaluation methods: (i) a task-specific evaluation function and (ii) a commercial LLM evaluation system. These methods operate independently and can be chosen by users, with the first focusing on predefined tasks and the second offering a scalable solution for adding tasks across various apps. And to mimic the real-world scenario, all tasks are evaluated by real-time states, which means all the contents are real-time, not predefined offline contents.
@@ 3.4.1 Evaluation Function
  • Element matching is the most commonly used evaluation method. It involves identifying key elements in an XML tree and comparing their attributes with ground-truth values. For instance, consider the task: “Open Downloads in Coursera and tell me how much storage is used.” In this case, the final state XML should contain an element that displays the total storage used by the app. The ground-truth value can be extracted by parsing the XML tree and then compared to the agent’s response. In more complex scenarios, multiple elements may need to be retrieved, and several conditions must be met to determine if the agent’s answer is correct. Additionally, when the XML data is insufficient, OCR (Optical Character Recognition) is used to extract text attributes directly from an element, serving as a substitute for XML parsing.
  • Action matching is used when evaluation requires verifying specific positions. For example, in the task: “Search for a flashlight on the Wish app, filter results by price under 100, then select the first item and add it to the wishlist,” the agent must click on the first item displayed in the search results. Action matching ensures that the click coordinates fall within the bounding box of the first item.
The primary limitation of function evaluation arises from the conditions and rules defined within the function itself. This is because it is challenging for coders to account for all possible corner cases and avoid imposing overly strict conditions on action history or state elements when determining whether a task has been successfully completed. For example, if an agent navigates to an incorrect page but eventually returns to the correct state after a few steps, it becomes difficult for coders to establish appropriate conditions to accurately handle such scenarios. Another example is that after completing a task, a pop-up window may appear and obscure critical elements on the screen, leading the evaluation function to misjudge the outcome.
@@ 3.4.2 LLM Evaluation System
Agent / Model Metric Easy Med. Hard Op. S. Q. M. Q. Overall V.S. Human
InternVL2-8B Func SR 21.3 1.3 0 14.0 2.0 0.0 10.5 94.0
LLM SR 22.3 5.2 0.0 16.1 4.1 0.0 12.4 94.5
ESAR 32.2 20.3 20.5 31.8 19.7 21.2 26.9 -
Qwen2.5-VL-7B Func SR 23.4 5.2 0.0 17.5 0.0 0.0 12.9 94.0
LLM SR 27.7 7.8 0.0 19.6 8.2 0.0 15.9 93.5
ESAR 39.8 27.2 29.4 36.4 25.6 28.3 33.2 -
UI-TARS-7B-SFT Func SR 28.7 9.1 0.0 21.0 0.0 0.0 16.9 93.0
LLM SR 34.0 15.6 0.0 23.8 16.3 0.0 21.9 93.0
ESAR 55.8 40.1 41.8 51.1 36.2 35.7 46.5 -
GPT-4o-2024-11-20 Func SR 5.3 0.0 0.0 3.5 2.0 0.0 2.5 96.0
LLM SR 8.5 0.0 0.0 4.9 2.0 0.0 3.9 95.5
ESAR 14.6 9.2 8.6 15.1 7.2 7.7 9.1 -
Claude-3.5-sonnet Func SR 11.7 2.6 0.0 8.4 2.0 0.0 6.5 94.0
LLM SR 13.8 2.6 0.0 9.8 2.0 0.0 8.8 93.9
ESAR 23.6 16.7 14.2 22.1 14.7 14.9 17.6 -
AppAgent Func SR 19.1 2.6 0.0 12.6 4.1 0.0 9.9 94.0
LLM SR 21.3 5.2 0.0 14.7 6.1 0.0 14.0 93.5
ESAR 26.7 20.5 18.6 27.1 21.7 16.8 23.8 -
Table 4: The evaluation results in A3. The top three rows are GUI-finetuned or pretrained agents. The bottom three rows are or utilize commercial generalist LLMs. “Func SR” stands for task success rate by evaluation function. “LLM SR” stands for task success rate by commvercial LLM evaluation. “EASR” stands for essential state achieved rate.
During the development of A3, we encountered similar challenges discussed above. Each task requires a unique evaluation function, which must be meticulously designed by coding experts with the ability to parse XML and define precise success conditions. However, even experienced developers may struggle to account for all possible scenarios. This reliance on manual coding and condition setting significantly slows down the creation of evaluation functions. To overcome these limitations, we propose a commercial LLM-based evaluation system that harnesses the advanced capabilities of large language models (LLMs), such as GPT, Gemini, and Claude, to enable fully autonomous task evaluation.
LLM evaluation system can be categorized into three methods, which are (i) final state evaluation, (ii) sequence state evaluation and (iii) essential state evaluation.
  • Final state evaluation is conducted by LLM evaluators, who assess whether a task has been successfully completed based solely on the final execution state. This approach enables quick and accurate evaluation for tasks where the final state provides all the necessary information for judgment, such as “Open the notification page of CNN.” In those cases, the final state fully encapsulates the task requirements. However, this method can result in misjudgments if the final state changes unexpectedly or fails to include the information needed for evaluation.
  • Sequence state evaluation is designed to overcome the limitations of final state evaluation by assessing task completeness using a combined sequence of all states throughout the execution. In this approach, all screenshots are merged into a single image, and the corresponding XMLs are serialized. This method provides accurate and efficient evaluation for simpler tasks where the sequence length is fewer than six states. However, for longer sequences, the combined image becomes overly compressed and flat, leading to potential confusion and hallucination during evaluation.
  • Essential state evaluation offers the most comprehensive approach among the three evaluation methods. In this method, each task is broken down into a series of “essential states.” For example, the task “Open Gmail and select the ’Draft’ folder” can be divided into two essential states: [“Gmail is opened,” “Draft folder is selected”]. A sliding window is then applied to the entire sequence, grouping a set of states and XMLs, similar to sequence state evaluation. However, instead of evaluating the entire sequence, evaluators determine which essential states are achieved within the sliding window. The window moves incrementally from the start to the final state, and evaluators collect all the essential states that have been achieved. If the set of achieved essential states matches the complete list of essential states, the task is confidently deemed successful. This method also introduces an additional evaluation metric, the Essential State Achieved Rate (ESAR), which is calculated as the number of achieved essential states divided by the total number of essential states. ESAR provides a more granular performance assessment for agents. For example, even if a task fails, an agent that achieves 2 out of 3 essential states performs better than one that achieves only 1 essential state. More detailed essential states generation is in Appendix A.3
Hallucination is an inherent limitation of LLMs, leading to occasional misjudgments even in the most advanced commercial models. However, evaluator voting can significantly reduce the rate of misjudgment. We conducted experiments using three of the most advanced commercial LLMs, comparing their evaluations against human judgment on 50 tasks. The accuracy results are presented in Table 3. When evaluator voting is applied, the accuracy of the judgments increases to a conservative estimate of over 92%, demonstrating a high level of confidence in the evaluation outcomes.